A shuttle is being launched off a planet that has no air. The planet has a radius of 4.72×106 m and a mass of 3.22×1024 kg. The shuttle is fired off the planet at a speed of 8.40×103 m/s. A small satellite is orbiting 780. m above the planet. What is the speed of the shuttle as it passes the satellite?
Answer #1

By energy conservation,

PE + KE at surface = PE + KE while crossing the satellite

- GMm/R + 0.5mu^2 = - GMm/r + 0.5mv^2

- GM/R + 0.5u^2 = - GM/r + 0.5v^2

- 6.67e-11*3.22e24/(4.72e6 ) + 0.5*8.4e3^2 = -6.67e-11*3.22e24/(4.72e6 +780) + 0.5v^2

v = 8399 m/s answer